example:

You are a Housekeeping Manager for a large mountain resort, and you are trying to calculate your Wage Cost per Occupied Room for the day. The following information holds true

-43 employees worked today.

-The total hours worked by the employees was 344 hours.

-The average hourly rate of employees is $12/hour.

-There were 927 rooms occupied today at the resort.

Q. What is your Wage Cost per Occupied Room today?
Answer #1

Total wage payment is (total hour worked * average rate per hour)= 344*12 = 4128

Wage cost / occupied room ( total wages paid / no. of occupied room) = is 4128 / 927 = 4.453